[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jun 18, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Jun 18 05:14:48 UTC 2016


HSU Shift Log for Sat Jun 18 2016 01:14:41 GMT-0400 (EDT)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: Thanks Mike for correcting minor plotting glitches.

Other Techincal issues: Loud tone/squeal occurred on Hotline from  
1830-1850Z; persistent hissing static sound continues and we probably would  
not be able to conduct a clear conference call due to high pitch of  
background noise. AT&T is aware of problem, but has done nothing to  
eliminate or reduce background noise the past 3 days.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160618/12d220bc/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list